If 1.31 g Ar are added to 2.21 atm He in a 2.00 L cylinder at 27.0 °C, what is the total pressure of the resulting gaseous mixture?
Answer #1

law of partialPYessures Doltons Accondiny t the mixtuve xeated by the total pr(SSuYe Su is epual to Jase edividual jaes of thmeles ot AY 1-319 O.03275 mols AY= 409/mas mols ot He в Пне PAY hHехRT V (0-03715 holes) x (o-0821 L-at mk(21+273)k 2 L PA= 0